OpenStudy (anonymous):

integrate absolute value of (6-x) dx from 5 to 9 using midpoint riemann

OpenStudy (anonymous):

\[\int\limits_{5}^{9}\left| 6-x \right|dx\]

OpenStudy (anonymous):

So the first thing I did was I split up the integral to be this: \[\int\limits_{5}^{6}6-x dx + \int\limits_{6}^{9}-6+xdx\]

OpenStudy (anonymous):

Is that the right first step?

OpenStudy (anonymous):

The midpoint riemann sum formula is: \[\Delta x \Sigma f(a+(j-\frac{ 1 }{ 2 })\Delta x)\]

OpenStudy (anonymous):

How many subintervals from 5 to 9?

OpenStudy (anonymous):

4 subintervals.